# Probabilistic Counters for Privacy Preserving Data Aggregation11footnotemark: 1

### v5

@@ 6 Formulation of main results
In this section we state main results. Proofs are postponed to Section 8 for the convenience of the reader.
@@ 6.1 Morris Counter Privacy
In this subsection we investigate Morris Counter in terms of $(\varepsilon,\delta)$-DP. Here we present the main result:
@@ 6.2 General result on Morris’ Counter privacy
where $c$ is some positive constant such that $\lceil c\log(\ln(n))\rceil\geqslant 1$, for big enough $n$.
@@ 6.3 MaxGeo Counter Privacy
In this subsection, we present a theorem that shows the privacy guarantees of MaxGeo Counter. Assume that we have $n$ increment requests. In the case of MaxGeo Counter it means that we generate random variables $X_{1},\ldots,X_{n}$, where $X_{i}\sim\mathrm{Geo}({1}/{2})$ are pairwise independent. Ultimately the result of the counter is maximum over all $X_{i}$’s, namely $X=\max(X_{1},\ldots,X_{n})$. Now we are ready to present our second main contribution.
@@ 7.1 Discussion on arragements of the survey scenario
We usually consider probabilistic counters to be some kind of estimator. There is always a trade-off between its precision and the quality of privacy. Ultimately, it means that for fixed privacy parameters $(\varepsilon,\delta)$ we can calculate the minimum number of increment requests necessary to satisfy given privacy parameters. This can be done by artificially adding them before actually collecting data. Of course, it has to be taken into account that the initially added value should be subtracted from the final estimation of the appropriate cardinality before publication, and this change can impact the precision of the estimation (especially when the expected number of increment requests is very small). If we can perform such a preprocessing, then for every $(\varepsilon,\delta)$, we can easily know how many artificial counts have to be added. Nevertheless, we have to be aware that if the total number of increment requests is small, we may obtain poor approximation, so this approach should be used whenever the privacy is much more important than the precision.
Note that, in light of our theorems 1 and 3, both Morris Counter and MaxGeo Counter preserve differential privacy in such a scenario. Assume at least $n$ users holding ${}^{\prime}1^{\prime}$, therefore at least $n$ incrementation requests. See that we can either know it based on domain knowledge (e.g., we expect that at least some fraction of users will send ${}^{\prime}1^{\prime}$ based on similar surveys) or add $x$ counts to the counter artificially initially. The number $x$ should be chosen according to the maximal amenable value of $\varepsilon$ parameter for a given application, but we recommend choosing rather small values of $x$. Obviously, in the case of artificial counts, it has to be taken into account when estimating the final sum. Using Morris Counter we obtain $\left(L(n),0.00033\right)$-DP with
@@
Consider a result of Morris Counter with a small number $n$ of increment requests (for example, a number of respondents suffering from a rare illness). Therefore, we will likely require the $\varepsilon$ parameter to be at most some threshold, e.g. $1$. Therefore, from Theorem 1, we should add $x$ counts where $L(x)\leqslant{16}/{(x-8)}\leqslant 1$, so $x\geqslant 24$. Note that we do not include $n$ in the above formula since it is not known in advance. Therefore, using Morris Counter, the above survey is at least $(L(n+24),0.00033)$-DP. However, the estimator should be modified as well, i.e., $M^{\prime}_{n}=M_{n+24}$, so $\hat{n^{\prime}}=\max\{2^{M^{\prime}_{n}}-26,0\}$ (since $2^{M^{\prime}_{n}}$ may be smaller than $26$).
@@
Assume that we have at least $n=200$ increment requests. From Theorem 1, we have $L(n)\leqslant{16}/{(n-8)}\leqslant 0.08334$. Therefore, using Morris Counter, the above survey is $(0.08334,0.00033)$-DP.
On the other hand, using MaxGeo Counter for a given $\varepsilon$ and $\delta$ we get $\left(\varepsilon,\delta\right)$-DP as long as $n\geqslant\dfrac{\ln(\delta)}{\ln\left(1-2^{-l_{\varepsilon}}\right)}$, where $l_{\varepsilon}=\left\lceil\log\left(1+{1}/{\varepsilon}\right)\right\rceil$. Here we present an example.
@@
Let $\varepsilon=0.5$ and $\delta={1}/{D^{2}}$, where $D$ is the number of all survey participants. After using our theorem and straightforward calculations, we have $n\geqslant 7\ln(D)~{}.$ Say we will have $\lfloor\exp(20)\rfloor$ participants. Then if we have at least $140$ increment requests, we satisfy $(0.5,{1}/{D^{2}})$-DP.
Note that from a differential privacy perspective, both the general PCSA algorithm and HyperLogLog can be seen as arbitrary postprocessing performed on the $m$ MaxGeo counters. Moreover, as each response goes to one counter only, they are independent of each other, so we can use the parallel composition theorem (see [5]).
@@ 7.2 Comparison of Morris and MaxGeo Counters
In this subsection, we compare data aggregation algorithms’ privacy and storage properties based on one of the investigated counters or the standard Laplace method.
We start with auxiliary remarks for the privacy of MaxGeo Counter. For instance, see that if $\delta$ and $n$ are fixed, then from Theorem 3 and $l_{\varepsilon}\leqslant\left\lceil\ln\left(1+\varepsilon^{-1}\right)\right\rceil$ we obtain that
@@
We want to optimize $\varepsilon(n)$, so we are going to consider $\varepsilon_{0}(n)$ defined as the right-hand side of (8). In order to limit it let us consider the following function of $x\in\mathbf{R}_{+}$:
@@
Naturally, then $\varepsilon_{0}(n)\geqslant\psi(n,\delta)=-{\ln(\delta)}/{n}+O(n^{-2})$. Since $\psi$ is decreasing with respect to $x$, we will consider when $\varepsilon_{0}(n)$ changes. More precisely, consider a minimal $k$ such that $\varepsilon_{0}(n)<\psi(n-k,\delta)\leqslant\varepsilon_{0}(n-k)$, which occurs to be the neat upper bound for $\varepsilon(n)$. However, since $\varepsilon_{0}(n)$ is the non-ascending step function, we realize that
@@
If we denote $\phi(n,\delta):=\left(2^{\left\lfloor-\log\left(1-\delta^{\frac{1}{n}}\right)\right\rfloor-1}-1\right)^{-1}$, then we can summarize our recent considerations in a short way by $\psi(n,\delta)\leqslant\varepsilon(n)<\phi(n,\delta)$. Thence, in the case where we fix the parameter $\delta=0.00033$, we obtain
@@
Therefore, Morris and MaxGeo Counters behave quite similarly under comparable conditions, and Figure 3 confirms this observation. Indeed, in Figure 3 we may see that the difference between the values of $\varepsilon(n)$ parameters for both counters decreases as $n$ increases.
@@
Realize that the previous conclusions remain true if $\delta(n)$ is not constant. This short observation enables us to obtain a more general result:
@@
Notice that in this case, both sequences of parameters tend to 00, which may be used as an advantage in applications, especially when we expect that the total number of increment requests will be very large. However, we emphasize that this requires $\delta(n)$ to be negligible.
@@
Estimator $\hat{n}$ $n+\mathcal{L}(n/16)$ $22^{M_{n}}-$ 2 $\sim 0.89\cdot 2^{\frac{1}{2}\sum_{i=1}^{2}M_{n}[i]}$
$\mathrm{Var}(\hat{n})$ $\frac{n^{2}}{128}$ $\frac{n^{2}+n}{2}$ $\sim 0.845\cdot n^{2}$
@@
In Figure 4, we may briefly see that probabilistic counters can be used for data aggregation in order to decrease the memory usage in exchange for a slight increase of $\delta$ parameter of differential privacy and wider confidence intervals (lower accuracy). In recent years Big Data related problems became very popular. Note that this kind of application makes major use of memory. When the server aggregates many different data, standard solutions may cause a serious problem with data storage, which can be encountered by using the idea based on a probabilistic counter instead.
@@
For every question, we would like to estimate the number of people who answered yes, but we want to guarantee differential privacy property at a reasonable level. Realize that if the number of yes answers is very small for some questions (e.g., when the question is about a very rare disease), then the number of no answers may be counted instead. Obviously, this method gains privacy, but loses precision.
According to Figure 4, if we use the Laplace method, then we may need approximately $100\log(10^{8})=2657.54\ldots$ bits to store the counters. However, if we use Morris Counter instead, about $100\log(\log(10^{8}))=473.20\ldots$ bits are needed. Let us note that all terms $O(1)$ in the ”Average memory” row of Figure 4 are bounded by $1$. Hence their impact is negligible from a practical point of view.
One may also complain about the heavy use of Pseudo Random Number Generator that probabilistic counters make. However, this problem may be resolved by generating the number of increment requests, which have to be forgotten until the next update of the counter by utilizing appropriate geometric distributions (see, for instance, [46] for a similar approach applied to reservoir sampling algorithm). This way, the use of PRNG can be substantially reduced.
@@ 8.1 Main Theorem for Morris Counter (Theorem 1)
The proof is complicated and very technical. In order to better understand it, we are going to provide a presentation of a plan and main ideas beneath the parts of the proof. Let $n_{k}:=2^{k}+1$ for $k\in\mathbb{N}$. We introduce $\mathcal{P}_{k}^{(c)}$ to be the probability of the Morris Counter $M_{n}$ being $k+c$ after $n_{k}$ increment requests, i.e. $p_{n_{k},k+c}$. The ”special” sequences $(\mathcal{P}_{k}^{(c)})_{k}$ play a crucial role in the proof.
@@
During the first phase, we consider a concentration of Morris Counter in the vicinity of its mean value. More precisely, we show that Morris Counter after $n$ increment requests takes values in relatively small intervals $I_{n}$ with probability $1-\delta$ (note that then $M_{n}$ satisfies the condition (2) for $S_{n}=I_{n}$), where $I_{n}$ is defined as in (4) and $\delta$ is some small constant, which arises from the proof. Note that $I_{n}$ may be interpreted as confidence intervals at level $1-\delta$ (see e.g. [47]). This phase is divided into lemmas 1 and 2. The first one uniformly bounds the formula for probabilities given by the Theorem below, due to Flajolet:
@@
In Lemma 1, we sum up bounds on $p_{n,l}$ to obtain small upper bound for $\delta_{1}:=\mathbb{P}\left(M_{n}\leqslant\left\lceil\log(n)\right\rceil-5\right)$. The same bounds cannot be efficiently utilized in the proof of Lemma 2. Instead, it couples $M_{n}$ with a process $X_{n}$, which increases during first $\left\lceil\log(n)\right\rceil+1$ steps, and then follow the same rule of update, so $M_{n}\leqslant X_{n}$ almost surely. Therefore $\delta_{2}:=\mathbb{P}\left(M_{n}\geqslant\left\lceil\log(n)\right\rceil+5\right)\leqslant\mathbb{P}\left(X_{n}\geqslant\left\lceil\log(n)\right\rceil+5\right)$, which is much easier to bound from definition. Let us note that such coupling cannot be used in the proof of Lemma 1. The first phase is summarized by Theorem 5, i.e., establishes $\delta=\delta_{1}+\delta_{2}$.
In the second phase we show that $(\mathcal{P}_{k}^{(4)})_{k}$ is descending for big enough $k$ and $(\mathcal{P}_{k}^{(5)})_{k}$ is ascending for big enough $k$ (Lemma 6). A change of monotonicity is the first obstacle in the proof of the main theorem. The main idea beneath the proof of this fact is to calculate the differences between consecutive elements of the considered ”special” sequences by representing them as the sums via application of Flajolet’s Theorem 4 and realizing that usually at most first ten terms of the sums are crucial (on the other hand, taking less than eight terms is rarely sufficient). This is the second issue, which makes the proof so complicated. Let us notice that Theorem 4 presents an explicit formula for $\mathbb{P}\left(M_{n}=l\right)$, which (as we may experience in A) is not convenient to analyze. However, it is simple enough to find the values numerically (also note that recursive Definition 3 provides those probabilities easily as well. However, this approach is inefficient in terms of memory and time for a big number of requests $n$). Therefore, by precise analysis, we can finally check some sums numerically and obtain the thesis of Lemma 6 for $k\geqslant 15$. However, numerically, one can extrapolate it to some smaller $k$ as well (remark that this proof does not work for small values of $k$, since not only ten terms of the aforementioned sums are important). Lemma 6 can be directly used to show that $\mathcal{P}_{k}^{(4)}\leqslant 2^{7}\mathcal{P}_{k}^{(5)}$ for $k\geqslant 7$ (Claim 1). Let us note that in fact, this result is not true for $k<7$.
@@
The fourth phase (i.e. Lemma 3) shows that if $p_{n_{k},k+c}\leqslant 2^{c+3}p_{n_{k},k+c+1}$, for $c\in[-k:4]$, then the same is true if we substitute $n_{k}$ with a bigger number (i.e., this property is increasing with respect to the $n$ parameter). In order to apply this result, we have to satisfy some starting conditions. We have numerically checked the appropriate condition for $k=7$ (presented later, in Table 1). Therefore, the second phase of the proof of Theorem 1 justifies the assumptions of Lemma 3 with respect to the $k$ parameter, as long as $k\geqslant 7$ and the third phase let us obtain the appropriate assumptions with respect to $c$. One can check that for $k<7$, an analogous assumption is not true.
The latter phase begins with the application of the result from the previous part. We obtain that $\varepsilon$ is at most $L(n)$ (provided in the formulation of Theorem 1), for $k\geqslant 7$. The last piece of this puzzle is justified by a numerical evaluation for $k<7$ (presented later, in Figure 5), which ends the $\varepsilon$ phase, and so the whole proof.
@@
Let us begin with a reminder. First of all, $M_{n}$ is ranged in $\mathbb{N}_{0}$ and, moreover, $I_{n}\subseteq\left[\left\lceil\log(n)\right\rceil-4:\left\lceil\log(n)\right\rceil+4\right]$. We provide few facts about the concentration of the distribution of the random variable $M_{n}$, or more precisely about the probability that $M_{n}$ will be outside the interval $I_{n}$.
@@
Let $M_{n}$ be the state of the Morris counter after $n$ increment requests. Then
@@
An increasing sequence $\prod\limits_{i=1}^{k}\left(1-2^{-i}\right)^{-1}$ that arose in Theorem 4 will be indicated by $r_{k}$ (with $r_{0}=1$) and we denote its limit $\prod\limits_{i=1}^{\infty}\left(1-2^{-i}\right)^{-1}=3.46274\ldots$ by $R$.
Let us mention that with the notions $q_{k}$ and $Q$ from [2], $r_{k}=\frac{1}{q_{k}}$ and $R=\frac{1}{Q}$.
@@
At first, we want to bound a lower tail of the distribution $\delta_{1}$. Here we would like to find a sufficient upper limit for the above probability. Assume that $l\leqslant\left\lceil\log(n)\right\rceil-5$. Let us realize that $r_{k}\leqslant R$ and that $y=-x+1$ is a tangent line to the plot of $y=-x(x-1)$ in the point $(x,y)=(1,0)$. Therefore:
@@
Above formula will help us to limit the left tail od the distribution of $M_{n}$:
@@
The bound for $p_{n,l}$ obtained above, is useless when $l\geqslant\log(n)-2$, so it cannot be used in the next lemma for an symmetrical upper tail.
@@
Let $M_{n}$ be the state of the Morris Counter after $n$ increment requests. Then
@@
for $k\geqslant\left\lceil\log(n)\right\rceil+1$. Naturally, for $k\leqslant\left\lceil\log(n)\right\rceil+1$, we have $X_{k}\geqslant M_{k}$, so we may couple realizations of these two processes in such a way that whenever $X$ is incremented, then so is $M$ and if $M$ does not change, then $X$ does not increase as well (note that $X$ has at most the same probability of a positive increment as $M$ at any point of time).
To abbreviate the expressions let us denote $m=n-\left\lceil\log(n)\right\rceil-1$ and
@@
Realize that $\sum\limits_{k=3}^{m}k=(m-2)(m+3)/2$ and $\sum\limits_{k=3}^{m}k^{2}=(m-2)(2m^{2}+7m+15)/6$, so
@@
Note that when $m<3$ (that is, when $n<7$), then the above sums are empty, but on the other hand $\lceil\log(n)\rceil+5>n+1$, so the inequality is trivially true.$\qed$
@@
Realize that $\mathbb{P}\left(M_{n}\in[1:n+1]\right)=1$. This observation, together with lemmas 1 and 2 yield
@@
In this part of the investigation, we try to establish $\varepsilon(n)$ parameter of DP of $M_{n}$. In fact it remains to examine the property (1) in the interval $I_{n}$, as Theorem 5 entails (2) for $S_{n}=I_{n}$. Therefore, we are interested in finding the upper bound for maximal privacy loss for any $n\in\mathbb{N}$ and $k\in I_{n}$, namely:

### v6

@@ 6 Formulation of main results
In this section, we state main results. The proofs are postponed to Section 8 for the convenience of the reader.
@@ 6.1 Morris Counter Privacy
In this subsection, we investigate the Morris Counter in terms of $(\varepsilon,\delta)$-DP. Here we present the main result:
@@ 6.2 General result on Morris’ Counter privacy
where $c$ is some positive constant such that $\lceil c\log(\ln(n))\rceil\geqslant 1$, for large enough $n$.
@@ 6.3 MaxGeo Counter Privacy
In this subsection, we present a theorem that shows the privacy guarantees of MaxGeo Counter. Assume that we have $n$ increment requests. In the case of MaxGeo Counter, it means that we generate random variables $X_{1},\ldots,X_{n}$, where $X_{i}\sim\mathrm{Geo}({1}/{2})$ are pairwise independent. Ultimately, the result of the counter is maximum over all $X_{i}$’s, namely $X=\max(X_{1},\ldots,X_{n})$. Now we are ready to present our second main contribution.
@@ 7.1 Discussion on arragements of the survey scenario
We usually consider probabilistic counters to be some kind of estimator. There is always a trade-off between its precision and the quality of privacy. Ultimately, it means that for fixed privacy parameters $(\varepsilon,\delta)$ we can calculate the minimum number of increment requests necessary to satisfy given privacy parameters. This can be done by artificially adding them before actually collecting data. Of course, it has to be taken into account that the initial added value should be subtracted from the final estimation of the appropriate cardinality before publication, and this change can impact the precision of the estimation (especially when the expected number of increment requests is very small). If we can perform such a preprocessing, then for every $(\varepsilon,\delta)$, we can easily know how many artificial counts have to be added. Nevertheless, we have to be aware that if the total number of increment requests is small, we may obtain a poor approximation, so this approach should be used whenever the privacy is much more important than the precision.
Note that, in light of our theorems 1 and 3, both the Morris Counter and the MaxGeo Counter preserve differential privacy in such a scenario. Assume that at least $n$ users have $^{\prime}1^{\prime}$, therefore at least $n$ incrementation requests. See that we can either know it based on domain knowledge (e.g., we expect that at least some fraction of users will send $^{\prime}1^{\prime}$ based on similar surveys) or add $x$ counts to the counter artificially initially. The number $x$ should be chosen according to the maximal amenable value of the parameter $\varepsilon$ for a given application, but we recommend choosing rather small values of $x$. Obviously, in the case of artificial counts, this has to be taken into account when estimating the final sum. Using the Morris counter, we obtain $\left(L(n),0.00033\right)$-DP with
@@
Consider a result of Morris Counter with a small number $n$ of increment requests (for example, a number of respondents suffering from a rare illness). Therefore, we will likely require the parameter $\varepsilon$ to be at most some threshold, e.g. $1$. Therefore, from Theorem 1, we should add $x$ counts where $L(x)\leqslant{16}/{(x-8)}\leqslant 1$, so $x\geqslant 24$. Note that we do not include $n$ in the above formula since it is not known in advance. Therefore, using the Morris Counter, the above survey is at least $(L(n+24),0.00033)$-DP. However, the estimator should be modified as well, i.e., $M^{\prime}_{n}=M_{n+24}$, so $\hat{n^{\prime}}=\max\{2^{M^{\prime}_{n}}-26,0\}$ (since $2^{M^{\prime}_{n}}$ may be smaller than $26$). margin:
@@
Assume that we have at least $n=200$ increment requests. From Theorem 1, we have $L(n)\leqslant{16}/{(n-8)}\leqslant 0.08334$. Therefore, using the Morris Counter, the above survey is $(0.08334,0.00033)$-DP.
On the other hand, using the MaxGeo Counter for a given $\varepsilon$ and $\delta$ we get $\left(\varepsilon,\delta\right)$-DP as long as $n\geqslant\dfrac{\ln(\delta)}{\ln\left(1-2^{-l_{\varepsilon}}\right)}$, where $l_{\varepsilon}=\left\lceil\log\left(1+{1}/{\varepsilon}\right)\right\rceil$. Here we present an example.
@@
Let $\varepsilon=0.5$ and $\delta={1}/{D^{2}}$, where $D$ is the number of all the survey participants. After using our theorem and straightforward calculations, we have $n\geqslant 7\ln(D)~.$ Say we will have $\lfloor\exp(20)\rfloor$ participants. Then if we have at least $140$ increment requests, we satisfy $(0.5,{1}/{D^{2}})$-DP.
Note that from a differential privacy perspective, both the general PCSA algorithm and HyperLogLog can be seen as arbitrary postprocessing performed on the $m$ MaxGeo counters. Moreover, since each response goes to one counter only, they are independent of each other, so we can use the parallel composition theorem (see [18]).
@@ 7.2 Comparison of Morris and MaxGeo Counters
In this subsection, we compare the’ privacy and storage properties of data aggregation algorithms based on one of the investigated counters or the standard Laplace method.
We start with auxiliary remarks for the privacy of MaxGeo Counter. For instance, see that if $\delta$ and $n$ are fixed, then from Theorem 3 and $l_{\varepsilon}\leqslant\left\lceil\ln\left(1+\varepsilon^{-1}\right)\right\rceil$ we obtain
@@
We want to optimise $\varepsilon(n)$, so we will consider $\varepsilon_{0}(n)$ defined as the right-hand side of (8). In order to limit it let us consider the following function of $x\in\mathbf{R}_{+}$:
@@
Naturally, then $\varepsilon_{0}(n)\geqslant\psi(n,\delta)=-{\ln(\delta)}/{n}+O(n^{-2})$. Since $\psi$ is decreasing with respect to $x$, we will consider when $\varepsilon_{0}(n)$ changes. More precisely, consider a minimal $k$ such that $\varepsilon_{0}(n)<\psi(n-k,\delta)\leqslant\varepsilon_{0}(n-k)$, which appears to be the neat upper bound for $\varepsilon(n)$. However, since $\varepsilon_{0}(n)$ is the non-ascending step function, we realise that
@@
If we denote $\phi(n,\delta):=\left(2^{\left\lfloor-\log\left(1-\delta^{\frac{1}{n}}\right)\right\rfloor-1}-1\right)^{-1}$, then we can summarise our recent considerations in a short way by $\psi(n,\delta)\leqslant\varepsilon(n)<\phi(n,\delta)$. Thence, in the case where we fix the parameter $\delta=0.00033$, we obtain
@@
Therefore, Morris and MaxGeo Counters behave quite similarly under comparable conditions, and Figure 3 confirms this observation. Indeed, in Figure 3 we can see that the difference between the values of the parameters $\varepsilon(n)$ for both counters decreases as $n$ increases.
@@
Realise that the previous conclusions remain true if $\delta(n)$ is not constant. This short observation enables us to obtain a more general result:
@@
Notice that in this case, both sequences of parameters tend to 0, which may be used as an advantage in applications, especially when we expect that the total number of increment requests will be very large. However, we emphasise that this requires that $\delta(n)$ be negligible.
@@
Estimator $\hat{n}$ $n+\mathcal{L}(n/16)$ $2^{M_{n}}-2$ $\sim 0.89\cdot 2^{\frac{1}{2}\sum_{i=1}^{2}M_{n}[i]}$
[…]
@@
In Figure 4, we may briefly see that probabilistic counters can be used for data aggregation to decrease memory usage in exchange for a slight increase of the parameter $\delta$ of differential privacy and wider confidence intervals (lower accuracy). In recent years Big Data related problems became very popular. Note that this kind of application makes major use of memory. When the server aggregates many different data, standard solutions may cause a serious problem with data storage, which can be encountered by using the idea based on a probabilistic counter instead.
@@
For every question, we would like to estimate the number of people who answered yes, but we want to guarantee the differential privacy property at a reasonable level. Realise that if the number of yes answers is very small for some questions (e.g., when the question is about a very rare disease), then the number of no answers may be counted instead. Obviously, this method gains privacy, but loses precision.
According to Figure 4, if we use the Laplace method, then we may need approximately $100\log(10^{8})=2657.54\ldots$ bits to store the counters. However, if we use the Morris Counter instead, about $100\log(\log(10^{8}))=473.20\ldots$ bits are needed. Note that all terms $O(1)$ in the ”Average memory” row of Figure 4 are bounded by $1$. Hence, its impact is negligible from a practical point of view.
One may also complain about the heavy use of the pseudo random number generator (PRNG) that probabilistic counters make. However, this problem may be resolved by generating the number of increment requests which have to be forgotten until the next update of the counter by using appropriate geometric distributions (see, for example, [33] for a similar approach applied to the reservoir sampling algorithm). This way, the use of PRNG can be substantially reduced.
@@ 8.1 Main Theorem for Morris Counter (Theorem 1)
The proof is complicated and very technical. In order to better understand it, we are going to provide a presentation of a plan and main ideas beneath the parts of the proof. Let $n_{k}:=2^{k}+1$ for $k\in\mathbb{N}$.
We introduce $\mathcal{P}_{k}^{(c)}$ as the probability that the Morris Counter $M_{n}$ being $k+c$ after $n_{k}$ increment requests, i.e., $p_{n_{k},k+c}$. The ”special” sequences $(\mathcal{P}_{k}^{(c)})_{k}$ play a crucial role in the proof.
@@
During the first phase, we consider a concentration of Morris Counter in the vicinity of its mean value. More precisely, we show that the Morris Counter after $n$ increment requests takes values in relatively small intervals $I_{n}$ with probability $1-\delta$ (note that then $M_{n}$ satisfies condition (2) for $S_{n}=I_{n}$), where $I_{n}$ is defined as in (4) and $\delta$ is some small constant, which arises from the proof. Note that $I_{n}$ can be interpreted as confidence intervals at level $1-\delta$ (see, e.g. [40]). This phase is divided into lemmas 1 and 2. The first one uniformly bounds the formula for probabilities given by the Theorem below, due to Flajolet:
@@
In Lemma 1, we sum up the bounds on $p_{n,l}$ to obtain a small upper bound for $\delta_{1}:=\mathbb{P}\left(M_{n}\leqslant\left\lceil\log(n)\right\rceil-5\right)$. The same bounds cannot be utilised efficiently in the proof of Lemma 2. Instead, it couples $M_{n}$ with a process $X_{n}$, which increases during the first $\left\lceil\log(n)\right\rceil+1$ steps, and then follows the same update rule, so $M_{n}\leqslant X_{n}$ almost surely.
Therefore $\delta_{2}:=\mathbb{P}\left(M_{n}\geqslant\left\lceil\log(n)\right\rceil+5\right)\leqslant\mathbb{P}\left(X_{n}\geqslant\left\lceil\log(n)\right\rceil+5\right)$, which is much easier to bound from definition. Note that such a coupling cannot be used in the proof of Lemma 1. The first phase is summarised by Theorem 5, i.e., establishes $\delta=\delta_{1}+\delta_{2}$.
In the second phase, we show that $(\mathcal{P}_{k}^{(4)})_{k}$ is descending for large enough $k$ and $(\mathcal{P}_{k}^{(5)})_{k}$ is ascending for big enough $k$ (Lemma 6). A change of monotonicity is the first obstacle in the proof of the main theorem. The main idea beneath the proof of this fact is to calculate the differences between consecutive elements of the considered ”special” sequences by representing them as the sums via application of Flajolet’s Theorem 4 and realising that usually at most first ten terms of the sums are crucial (on the other hand, taking less than eight terms is rarely sufficient). This is the second issue, which makes the proof so complicated. Let us note that Theorem 4 presents an explicit formula for $\mathbb{P}\left(M_{n}=l\right)$, which (as we may experience in A) is not convenient to analyse. However, it is simple enough to find the values numerically (also note that recursive Definition 3 provides those probabilities easily as well, However, this approach is inefficient in terms of memory and time for a large number of requests $n$). Therefore, by precise analysis, we can finally check some sums numerically and obtain the thesis of Lemma 6 for $k\geqslant 15$. However, numerically, one can extrapolate it to some smaller $k$ as well (remark that this proof does not work for small values of $k$, since not only ten terms of the aforementioned sums are important). The lemma 6 can be used directly to show that $\mathcal{P}_{k}^{(4)}\leqslant 2^{7}\mathcal{P}_{k}^{(5)}$ for $k\geqslant 7$ (Claim 1). Let us note that, in fact, this result is not true for $k<7$.
@@
The fourth phase (i.e. Lemma 3) shows that if $p_{n_{k},k+c}\leqslant 2^{c+3}p_{n_{k},k+c+1}$, for $c\in[-k:4]$, then the same is true if we substitute $n_{k}$ with a bigger number (i.e., this property is increasing with respect to the $n$ parameter). In order to apply this result, we have to satisfy some starting conditions. We have numerically checked the appropriate condition for $k=7$ (presented later, in Table 1). Therefore, the second phase of the proof of Theorem 1 justifies the assumptions of Lemma 3 with respect to the parameter $k$, as long as $k\geqslant 7$ and the third phase let us obtain the appropriate assumptions with respect to $c$. One can check that for $k<7$, an analogous assumption is not true.
The latter phase begins with the application of the result from the previous part. We obtain $\varepsilon$ is at most $L(n)$ (provided in the formulation of Theorem 1), for $k\geqslant 7$. The last piece of this puzzle is justified by a numerical evaluation for $k<7$ (presented later, in Figure 5), which ends the $\varepsilon$ phase and so the whole proof.
@@
Let us begin with a reminder. First, $M_{n}$ is ranged in $\mathbb{N}_{0}$ and, moreover, in $I_{n}\subseteq\left[\left\lceil\log(n)\right\rceil-4:\left\lceil\log(n)\right\rceil+4\right]$. We provide few facts about the concentration of the distribution of the random variable $M_{n}$, or more precisely about the probability that $M_{n}$ will be outside the interval $I_{n}$.
@@
Let $M_{n}$ be the state of the Morris counter after the $n$ increment requests. Then
@@
An increasing sequence $\prod\limits_{i=1}^{k}\left(1-2^{-i}\right)^{-1}$ that emerged in Theorem 4 will be indicated by $r_{k}$ (with $r_{0}=1$) and we denote its limit $\prod\limits_{i=1}^{\infty}\left(1-2^{-i}\right)^{-1}=3.46274\ldots$ by $R$.
Let us mention this with the notions $q_{k}$ and $Q$ from [24], $r_{k}=\frac{1}{q_{k}}$ and $R=\frac{1}{Q}$.
@@
At first, we want to bound a lower tail of the distribution $\delta_{1}$.
Here we would like to find a sufficient upper limit for the above probability. Assume that $l\leqslant\left\lceil\log(n)\right\rceil-5$.
Realise that $r_{k}\leqslant R$ and that y=x+1y=-x+1 is a tangent line to the plot of y=x(x1)y=-x(x-1) in the point $(x,y)=(1,0)$. Therefore:
@@
The above formula will help us limit the left tail of the distribution of $M_{n}$:
@@
The bound for $p_{n,l}$ obtained above is useless when $l\geqslant\log(n)-2$, so it cannot be used in the next lemma for a symmetric upper tail.
@@
Let $M_{n}$ be the state of the Morris Counter after the $n$ increment requests. Then
@@
for $k\geqslant\left\lceil\log(n)\right\rceil+1$. Naturally, for $k\leqslant\left\lceil\log(n)\right\rceil+1$, we have $X_{k}\geqslant M_{k}$, so we may couple realisations of these two processes in such a way that whenever $X$ increases, so is $M$ and if $M$ does not change, then $X$ does not increase as well (note that $X$ has at most the same probability of a positive increase as $M$ at any point in time).
To abbreviate the expressions, let us denote $m=n-\left\lceil\log(n)\right\rceil-1$ and
@@
Realise that $\sum\limits_{k=3}^{m}k=(m-2)(m+3)/2$ and $\sum\limits_{k=3}^{m}k^{2}=(m-2)(2m^{2}+7m+15)/6$, so
@@
[…]
@@
Realise that $\mathbb{P}\left(M_{n}\in[1:n+1]\right)=1$. This observation, together with the lemmas 1 and 2 yields
@@
In this part of the investigation, we try to establish the $\varepsilon(n)$ parameter of DP of $M_{n}$. In fact, it remains to examine the property (1) in the interval $I_{n}$, as Theorem 5 entails (2) for $S_{n}=I_{n}$. Therefore, we are interested in finding the upper bound for maximal privacy loss for any $n\in\mathbb{N}$ and $k\in I_{n}$, namely: